Multiple-Product Breakeven

Parker Pottery produces a line of vases and a line of ceramic figurines. Each line uses the same equipment and labor; hence, there are no traceable fixed costs. Common fixed cost equals $30,000. Parker's accountant has begun to assess the profitability of the two lines and has gathered the following data for last year:

Vases Figurines
Price $40 $70
Variable cost 30 42
Contribution margin $10 $28
Number of units 1,000 500

Required:

1. Compute the number of vases and the number of figurines that must be sold for the company to break even.

Break-even vases units
Break-even figurines units

2. Parker Pottery is considering upgrading its factory to improve the quality of its products. The upgrade will add $5,260 per year to total fixed cost. If the upgrade is successful, the projected sales of vases will be 1,500, and figurine sales will increase to 1,000 units. What is the new break-even point in units for each of the products?

Break-even vases units
Break-even figurines units
Answer #1

Break-even point is where the company is able to recover its fixed costs. It is a state of no profit no loss. The formula and detailed working are explained below with step by step approach. Please feel free to ask any doubt.

Answer 1)

Normal Break even point (units) Fixed cost/ Contribution per unit Total units produced- 1000+500 1500 Sales mix = Vases (10001 500X100 and Figurines(5001 500X100 Vases 66.67% Figurines 33.33% Contribution mix Vases ($10X66.67%) + Figurines ($28X3 3.33%) - $6.67+$9.33 S 16.00 Break even point (units) S30,000/S16 Units= 1875 Sales mix for Break even sales Vases Figurines Sales mix ratio Break even units 66.67%) 1,250 33.33% 625 1875 units)

___________________________________________________________________________________________________

Answer 2) Upgraded Sales mix = Vases (1 5002500)X 100 and Figurines( 1 000/2 500)X100 Vases = 60% Figurines-4000 Vases ($10X60%) + Figurines ($28X40%) - S6+ $11.2 -S 17.20 Contribution mix Break even point (units)(S30,000+$S5260)S17.20 Units 2050 Sales mix for Break even sales Vases Figurines Sales mix ratio Break even units (2050 units) 60.00%) 40.00% 1,230 820